OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 5. Find y when x = 10.

OpenStudy (anonymous):

y = k/x ok?

OpenStudy (anonymous):

now 5 = k/2 so k= nd now y=k/10